# Weighted Fisher divergence for high-dimensional Gaussian variational inference

### v1

@@ 7.2 Generalized linear mixed models
for $i=1,\dots,n$, $j=1,\dots,n_{i}$. Here, $\beta\in\mathbb{R}^{p}$ is the fixed effect, $b_{i}\in\mathbb{R}^{r}$ is the random effect characterizing the $i$th subject, and $X_{ij}\in\mathbb{R}^{p}$ and $Z_{ij}\in\mathbb{R}^{r}$ are the covariates. We assume $b_{i}\sim{\text{N}}(0,G^{-1})$ and let $G=WW^{\top}$ be the Cholesky decomposition of precision matrix $G$, where $W$ is a lower triangular matrix with positive diagonal entries. For unconstrained optimization of $W$, we introduce $W^{*}$ such that $W_{ii}^{*}=\log(W_{ii})$ and $W_{ij}^{*}=W_{ij}$ if $i\neq j$, and let $\zeta=\text{vech}(W^{*})$. Normal priors, $\beta\sim{\text{N}}(0,\sigma_{0}^{2}I_{p})$ and $\zeta\sim N(0,\sigma_{0}^{2}I_{r(r+1)/2})$, where $\sigma_{0}^{2}=100$ are assigned. The global variables are $\theta_{G}=(\beta^{\top},\zeta^{\top})^{\top}$ and the local variables are $\theta_{L}=(b_{1}^{\top},\dots,b_{n}^{\top})^{\top}$. Focusing on GLMMs with canonical links and the one-parameter exponential family, the log joint density of the model can be written as
$\quad+\log p(\beta)+\log p(\zeta)$
$=\sum_{i,j}\{y_{ij}\eta_{ij}-A(\eta_{ij})\}+n\log|W|$
where $A(\cdot)$ is the log-partition function and $C$ is a constant independent of $\theta$. For instance, $A(x)=\log(1+e^{x})$ for Bernoulli-distributed binary responses and $A(x)=\exp(x)$ for Poisson-distributed count responses. Let $H=\nabla_{\theta}^{2}\log h(\theta)$ and $H_{\theta_{i},\theta_{j}}=\nabla_{\theta_{i},\theta_{j}}^{2}\log h(\theta)$. The Hessian,
$H=\begin{bmatrix}H_{b_{1},b_{1}}&\ldots&0&H_{b_{1},\theta_{G}}\\ \vdots&\ddots&\vdots&\vdots\\ 0&\dots&H_{b_{n},b_{n}}&H_{b_{n},\theta_{G}}\\ H_{\theta_{G},b_{1}}&\ldots&H_{\theta_{G},b_{n}}&H_{\theta_{G}}\\ \end{bmatrix},$
has a sparse structure similar to $\Omega$. Both $\nabla_{\theta}h(\theta)$ and $H$ are derived in the supplement.
First, consider the epilepsy data (Thall and Vail, 1990) from a clinical trial with $n=59$ patients, who were randomly assigned to a drug, progabide ($\text{Trt}=1$), or a placebo ($\text{Trt}=0$). The response is the number of seizures experienced by each patient during four follow-up periods. Covariates include logarithm of the patient’s age at baseline, which is centered by subtracting the mean (Age), logarithm of 1/4 the number of seizures prior to the trial (Base), visit number coded as --0.3, --0.1, 0.1, 0.3 (Visit), and an indicator of the 4th visit (V4). We consider Poisson mixed models with random intercepts and slopes (Breslow and Clayton, 1993),
@@
for $i=1,...,n$, $j=1,...,4$.
KLD FDr SDr FDb SDb BaM MCMC
MMD Epi I 6.37$\pm$0.19 0.04$\pm$0.02 -0.06$\pm$0.02 2.89$\pm$0.05 6.48$\pm$0.22 9.51$\pm$1.52 -
Epi II 6.34$\pm$0.13 0.04$\pm$0.03 0.06$\pm$0.03 4.65$\pm$0.05 5.97$\pm$0.17 6.41$\pm$0.22 -
Toenail 4.17$\pm$0.05 0.35$\pm$0.01 0.11$\pm$0.01 1.42$\pm$0.01 2.53$\pm$0.02 2.46$\pm$0.02 -
Polypharmacy 5.66$\pm$0.05 0.49$\pm$0.00 0.19$\pm$0.01 2.40$\pm$0.01 3.79$\pm$0.03 3.80$\pm$0.03 -
$\frac{|\mu-\mu^{*}|}{\sigma^{*}}$ Epi I 0.07$\pm$0.05 2.04$\pm$1.62 2.30$\pm$2.09 0.28$\pm$0.19 0.07$\pm$0.05 0.07$\pm$0.05 -
Epi II 0.09$\pm$0.07 2.14$\pm$1.79 2.55$\pm$2.40 0.19$\pm$0.16 0.09$\pm$0.08 0.09$\pm$0.07 -
Toenail 0.21$\pm$0.14 1.20$\pm$1.30 1.57$\pm$2.17 0.68$\pm$0.65 0.36$\pm$0.20 0.38$\pm$0.22 -
Polypharmacy 0.18$\pm$0.10 1.01$\pm$1.19 1.34$\pm$2.12 0.44$\pm$0.38 0.22$\pm$0.13 0.23$\pm$0.13 -
$\frac{\sigma}{\sigma^{*}}$ Epi I 0.95$\pm$0.06 0.76$\pm$0.26 0.72$\pm$0.34 0.81$\pm$0.21 0.94$\pm$0.05 0.99$\pm$0.02 -
Epi II 0.94$\pm$0.08 0.94$\pm$0.29 0.69$\pm$0.28 0.88$\pm$0.18 0.95$\pm$0.08 0.97$\pm$0.07 -
Toenail 0.87$\pm$0.05 0.35$\pm$0.14 0.10$\pm$0.05 0.66$\pm$0.13 0.76$\pm$0.11 0.75$\pm$0.12 -
Polypharmacy 0.94$\pm$0.03 0.39$\pm$0.12 0.13$\pm$0.05 0.80$\pm$0.09 0.86$\pm$0.07 0.86$\pm$0.07 -
time Epi I 2.3 (40) 16.7 (47) 4.3 (12) 6.6 (24) 8.9 (26) 0.8 (0.3) 95.9
Epi II 3.9 (37) 73.3 (60) 26.0 (21) 16.7 (25) 28.5 (31) 13.1 (1.55) 380.8
Toenail 2.8 (11) 117.4 (30) 27.7 (7) 93.7 (30) 35.0 (10) 32.5 (0.3) 171.1
Polypharmacy 5.1 (10) 346.1 (30) 94.7 (7) 285.8 (30) 131.8 (11) 112.5 (0.4) 506.2
Table 7: GLMMs: Estimates of mean and standard deviation for MMD, normalized absolute difference in mode and standard deviation ratio.
Runtime is in seconds and number of iterations (in thousands) is given in brackets.
Next, consider the toenail data (De Backer et al., 1998) from a clinical trial comparing two oral antifungal treatments for toenail infections. Each of 294 patients was evaluated for up to seven visits, resulting in a total of 1908 observations. Patients were randomized to receive 250 mg of terbinafine ($\text{Trt}=1$) or 200 mg of itraconazole ($\text{Trt}=0$) per day. The response variable is binary, with 0 indicating no or mild nail separation and 1 for moderate or severe separation. Visit times in months ($t$) are standardized to have mean 0 and variance 1. A logistic random intercept model is fitted to this data,
@@
[…]
@@
In this section, we use batch size $B=5$ for FDb and SDb. For BAM, $B=100$ for the epilepsy data and $B=1000$ for the toenail and polypharmacy data. A larger batch size is used in the latter cases as BAM is very prone to ill-conditioned updates and the convergence is very slow for smaller batch sizes. The maximum number of iterations for the epilepsy data is $60,000$, and reduced to $30,000$ for the toenail and polypharmacy data as they contain much more observations.
Figure 7: GLMMs: Average lower bound every 1000 iterations.
Fig 7 shows that SDb is among the fastest to converge among SGD methods, achieving a lower bound higher than KLD for Epi I, Epi II and polypharmacy, and comparable to KLD for toenail. While FDb converges rapidly for Epi I and Epi II, it fails to converge by the maximum iteration of 30,000 for toenail and polypharmacy. Overall, methods based on FD (FDr, FDb) seem to converge much more slowly than methods based on SD (SDr, SDb).
From the MMD results in Table 7, FDr and SDr have the lowest MMD. In contrast, FDb and SDb yield substantial improvements over their reparameterization trick-based counterparts, achieving much higher MMDs. Among SGD methods based on the weighted Fisher divergence, SDb has the highest MMD, even surpassing KLD for Epi I. BaM outperforms KLD for Epi I and Epi II, but its performance falters for toenail and polypharmacy, where the dimension of $\theta$ is higher, reflecting the challenges that BaM faces in high-dimensional cases.
KLD is able to capture the marginal posterior mode of each variable most accurately, with comparable performance from SDb and BaM. While BaM captured the marginal posterior variance most accurately for Epi I and Epi II, its performance falls behind KLD for the higher-dimensional toenail and polypharmacy data. SDr underestimates the marginal posterior variance most severely, which is reminiscent of the “variational collapse” problem it faces in the mean-field setting.
@@
BaM is able to converge using the least number of iterations, by leveraging on closed-form updates and larger batch sizes. However, the computation cost per iteration of BaM is much higher than SGD methods, which can exploit the sparse structure of the precision matrix. This issue becomes more apparent as the dimension of $\theta$ increases. Among SGD methods, the runtime of KLD is the shortest. Methods based on FD tend to require more iterations to converge than those based on SD, resulting in longer runtime. SDb is able to converge in approximately the same number of iterations as KLD, although each iteration takes longer.
Fig 8 compares the marginal densities estimated using MCMC with variational approximations from KLD, SDb, and BaM for some variables in Epi I and polypharmacy. For Epi I in the first row, all variational approaches match the MCMC results very closely except for $\zeta$, where SDb underestimated the marginal posterior variance more severely than BaM and KLD. For the higher dimensional polypharmacy data in the second row, KLD matches the MCMC results most closely, while SDb and BaM display a higher tendency to underestimate the marginal posterior variance although the mode was captured more accurately in some cases.
7.3 Stochastic volatility models
@@
where $\phi\in(0,1)$. To allow unconstrained updates, we apply the transformations, $\sigma=\exp(\alpha)$ and $\phi=\{\exp(-\psi)+1\}^{-1}$. The set of local variables are $\theta_{L}=(b_{1},\dots,b_{n})^{\top}$ and the set of global variables are $\theta_{G}=(\alpha,\lambda,\psi)^{\top}$. We consider the prior $\theta_{G}\sim{\text{N}}(0,\sigma_{0}^{2}I)$, where $\sigma_{0}^{2}=10$. For this model, $b_{i}$ is independent of $b_{j}$ given $\theta_{G}$ a posteriori if $|i-j|>1$ The log joint density is
where $C$ is a constant independent of $\theta$. The gradient $\nabla_{\theta}h(\theta)$ and the Hessian,
$\nabla_{\theta}^{2}\log h(\theta)=\begin{bmatrix}H_{b_{1},b_{1}}&H_{b_{1},b_{2}}&\ldots&0&H_{b_{1},\theta_{G}}\\ H_{b_{2},b_{1}}&H_{b_{2},b_{2}}&\ldots&0&H_{b_{2},\theta_{G}}\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\dots&H_{b_{n},b_{n}}&H_{b_{n},\theta_{G}}\\ H_{\theta_{G},b_{1}}&H_{\theta_{G},b_{2}}&\ldots&H_{\theta_{G},b_{n}}&H_{\theta_{G},\theta_{G}}\\ \end{bmatrix},$
which has the same sparsity structure as $\Omega$ in the variational approximation, are derived in the supplement.
KLD FDr SDr FDb SDb BaM MCMC
MMD GBP 3.83$\pm$0.09 0.45$\pm$0.01 1.03$\pm$0.01 0.49$\pm$0.01 3.26$\pm$0.08 3.59$\pm$0.09 -
DEM 5.34 $\pm$ 0.27 0.36 $\pm$ 0.01 0.78 $\pm$ 0.01 0.40 $\pm$ 0.01 6.02 $\pm$ 0.15 - -
$\frac{|\mu-\mu^{*}|}{\sigma^{*}}$ GBP 0.07$\pm$0.05 0.43$\pm$0.43 0.36$\pm$0.32 0.38$\pm$0.37 0.04$\pm$0.03 0.05$\pm$0.04 -
DEM 0.09$\pm$0.07 1.15$\pm$0.91 1.19$\pm$0.74 0.99$\pm$0.77 0.06$\pm$0.05 - -
time GBP 7.1 (27) 12.7 (30) 6.6 (15) 621.8 (30) 419.2 (16) 1521.5 (3.25) 3573.2
DEM 9.7 (19) 27.3 (30) 23.7 (25) 2748.0 (30) 1281.3 (11) - 5764.9
Table 8: SSMs: Estimates of mean and standard deviation for MMD and normalized absolute difference in mode.
Runtime is in seconds and number of iterations (in thousands) is given in brackets.
We analyze two datasets from Garch in the R package Ecdat. The first contains $n=945$ observations of the weekday exchange rates of the U.S. Dollar against the British Pound (GBP) from 1 Oct 1981 to 28 June 1985. The second contains $n=1866$ observations of the weekday exchange rates for the U.S. Dollar against the German Deutschemark (DEM) from 2 Jan 1980 to 21 June 1987. For both datasets, the mean-corrected log-return series $\{y_{t}\}$ is derived from exchange rates $\{r_{t}\}$ using
@@
We use $B=10$ for FDb and SDb, while BaM is executed with $B=50$ for GBP. We have tried larger batch sizes for BaM, but these resulted in severely ill-conditioned updates. For the higher-dimensional DEM, we cannot get BaM to converge despite trying various batch sizes. The challenge of inferring a full covariance matrix of order $d=1869$ for BaM here is immense, not to mention the computational cost of inverting matrices of this order. The maximum number of iterations is set as $30,000$.
From the MMD values in Table 8, FDr, SDr and FDb perform very poorly, compared to the rest. SDb yields the best performance among SGD methods optimizing the weighted Fisher divergence. Its MMD values are close to those of KLD for GBP, and even surpass KLD for the higher-dimensional DEM. SDb and BaM are also able to capture the mode more accurately than KLD for both datasets. In terms of runtime, KLD is the most efficient approach, attaining very high MMD values with the shortest runtime. While BaM requires fewer iterations to converge, the computational cost per iteration is much higher, leading to long runtimes. For instance, BaM requires 1521 seconds to converge for GBP, which is nearly triple the runtime of SDb. As the dimensionality increases, BaM also becomes more unstable and prone to ill-conditioned updates making convergence very difficult, while SDb provides a more reliable and computationally feasible approach for optimizing the batch approximated SD.
Figure 9: SSMs: Average lower bound over every 1000 iterations and boxplot of MMD values for different batch sizes for SDb.
Fig 9 presents the impact of varying the batch size for SDb in terms of the rate of convergence and quality of approximation measured by MMD. Increasing the batch size clearly leads to faster convergence and improved approximation quality. The total runtime as shown in the legends of the first row of Fig 9 is also reduced as lesser number of iterations is required for convergence. This suggests that larger batch sizes can enhance both the stability and accuracy of SDb.
Fig 10 shows the marginal posterior density estimates from MCMC, KLD and SDb ($B=10,100$) for some local variables and all the global variables from DEM. SDb can capture the marginal posterior mode more accurately than KLD, especially for each of the global variables, but has a higher tendency to underestimate the posterior variance. Increasing the batch size from 10 to 100 helps in reducing underestimation of the posterior variance.
Figure 10: Marginal density estimates for some local and global variables in DEM.

### v2

@@ 7.2 Generalized linear mixed model
for $i=1,\dots,n$, $j=1,\dots,n_{i}$. Here, $\beta\in\mathbb{R}^{p}$ is the fixed effect, $b_{i}\in\mathbb{R}^{r}$ is the random effect characterizing the $i$th subject, and $X_{ij}\in\mathbb{R}^{p}$ and $Z_{ij}\in\mathbb{R}^{r}$ are the covariates. We assume $b_{i}\sim{\text{N}}(0,G^{-1})$ and let $G=WW^{\top}$ be the Cholesky decomposition of precision matrix $G$, where $W$ is a lower triangular matrix with positive diagonal entries. For unconstrained optimization of $W$, we introduce $W^{*}$ such that $W_{ii}^{*}=\log(W_{ii})$ and $W_{ij}^{*}=W_{ij}$ if $i\neq j$, and let $\zeta=\text{vech}(W^{*})$. Normal priors, $\beta\sim{\text{N}}(0,\sigma_{0}^{2}I_{p})$ and $\zeta\sim{\text{N}}(0,\sigma_{0}^{2}I_{r(r+1)/2})$, where $\sigma_{0}^{2}=100$ are assigned. The global variables are $\theta_{G}=(\beta^{\top},\zeta^{\top})^{\top}$ and the local variables are $\theta_{L}=(b_{1}^{\top},\dots,b_{n}^{\top})^{\top}$. We focus on GLMMs with canonical link functions and responses from the one-parameter exponential family. The gradient $\nabla_{\theta}\log h(\theta)$ and Hessian $H=\nabla_{\theta}^{2}\log h(\theta)$, which has a sparse structure analogous to that of $\Omega$, are derived in the supplement.
First, consider the epilepsy data (Thall and Vail, 1990) from a clinical trial with $n=59$ patients, who were randomly assigned to a drug, progabide ($\text{Trt}=1$), or a placebo ($\text{Trt}=0$). The response is the number of seizures experienced by each patient during 4 follow-up periods. Covariates include logarithm of the patient’s age at baseline, which is centered by subtracting the mean (Age), logarithm of 1/4 the number of seizures prior to the trial (Base), visit number coded as -0.3, -0.1, 0.1, 0.3 (Visit), and an indicator of the 4th visit (V4). We consider Poisson mixed models with random intercepts and slopes (Breslow and Clayton, 1993),
@@
for i=1,,ni=1,...,n, j=1,,4j=1,...,4.
KLD FDr SDr FDb SDb (p)BaM
[…]
Toenail 0.21 $\pm$ 0.13 1.21 $\pm$ 1.30 1.58 $\pm$ 2.20 0.67 $\pm$ 0.66 0.35 $\pm$ 0.21 0.33 $\pm$ 0.15
[…]
Toenail 0.11 $\pm$ 0.07 1.46 $\pm$ 1.20 1.83 $\pm$ 2.14 0.78 $\pm$ 0.61 0.36 $\pm$ 0.23 0.28 $\pm$ 0.18
[…]
Toenail 0.88 $\pm$ 0.05 0.35 $\pm$ 0.14 0.10 $\pm$ 0.05 0.67 $\pm$ 0.14 0.76 $\pm$ 0.11 0.80 $\pm$ 0.11
[…]
time Epi I 2.3 (40) 16.7 (47) 4.3 (12) 6.6 (24) 11.6 (34) 1.0 (0.4)
Epi II 5.9 (52) 73.3 (60) 26.0 (21) 16.7 (25) 39.4 (42) 17.7 (2.3)
Toenail 2.8 (11) 117.4 (30) 27.7 (7) 93.7 (30) 35.0 (10) 11.3 (1.35)
Polypharmacy 5.1 (10) 346.1 (30) 94.7 (7) 285.8 (30) 139.2 (12) 36.7 (1.75)
Table 6: GLMM. Mean and standard deviation of normalized absolute difference in mode and mean and standard deviation ratio (best values highlighted in bold). Runtime is in seconds and number of iterations (in thousands) is given in brackets. BaM is used for Epi I and Epi II, and pBaM is used for Toenail and Polypharmacy.
Next, consider the toenail data (De Backer et al., 1998) from a clinical trial comparing two oral antifungal treatments for toenail infections. Each of 294 patients was evaluated for up to 7 visits, resulting in a total of 1908 observations. Patients were randomized to receive 250 mg of terbinafine ($\text{Trt}=1$) or 200 mg of itraconazole ($\text{Trt}=0$) per day. The response variable is binary, with 0 indicating no or mild nail separation and 1 for moderate or severe separation. Visit times in months ($t$) are standardized to have mean 0 and variance 1. A logistic random intercept model is fitted to this data,
@@
Lastly, we analyze the polypharmacy data (Hosmer et al., 2013) which contains 500 subjects, each observed for drug usage over 7 years, resulting in 3500 binary responses. Covariates include Gender (1 for males, 0 for females), Race (0 for whites, 1 otherwise), Age ($\log(\text{age}/10)$) and INPTMHV (0 if there are no inpatient mental health visits and 1 otherwise). The number of outpatient mental health visits (MHV) is coded as $\text{MHV1}=1$ if $1\leq\text{MHV}\leq 5$, $\text{MHV2}=1$ if $6\leq\text{MHV}\leq 14$, and $\text{MHV3}=1$ if $\text{MHV}\geq 15$. We consider a logistic random intercept model,
@@
We set $B=5$ for FDb and SDb. For BAM, $B=100$ for the epilepsy data. For the higher-dimensional toenail and polypharmacy data, BaM is prone to ill-conditioned updates and converges very slowly with smaller batchsizes. Hence, we use pBaM for these two datasets with $B=32$, as recommended by Modi et al. (2025). We set the rank $K=32$ for toenail data and $K=64$ for polypharmacy data, which is higher in dimension, so a larger $K$ is used. The maximum number of iterations for epilepsy data is $60,000$, which is reduced to $30,000$ for the larger toenail and polypharmacy data.
Figure 7: GLMM. Lower bound averaged over every 1000 iterations.
Fig 7 shows that SDb is among the fastest to converge among SGD methods, achieving a higher lower bound than KLD for Epi I, Epi II and polypharmacy, and comparable to KLD for toenail. FDb converges rapidly for Epi I and Epi II, but fails to converge by the maximum number of iterations for toenail and polypharmacy.
Figure 8: GLMM. Boxplots of $M^{*}$ values.
From Fig 8, FDr and SDr have the lowest $M^{*}$, while FDb and SDb yield substantial improvements over their reparameterization trick based counterparts. Among SGD methods based on the weighted Fisher divergence, SDb has the highest $M^{*}$, even surpassing KLD for Epi I and Epi II. BaM also outperforms KLD for Epi I and Epi II. While pBaM outperforms SDb for toenail and polypharmacy, it still falls short of KLD.
From Table 6, KLD is often able to capture the marginal posterior mean and mode most accurately, with comparable performance from SDb and (p)BaM. While BaM captured the marginal posterior variance most accurately for Epi I and Epi II, pBaM falls behind KLD for toenail and polypharmacy. SDr underestimates the marginal posterior variance most severely, which is reminiscent of the “variational collapse” problem it faces in the mean-field setting. (p)BaM is able to converge using the least number of iterations, by leveraging closed-form updates and larger batchsizes. However, the computation cost per iteration of (p)BaM is much higher than SGD methods, which can exploit the sparse structure of the precision matrix. This issue becomes more apparent as the dimension of $\theta$ increases. Among SGD methods, KLD is the fastest. Methods based on FD tend to require more iterations to converge than those based on SD, resulting in longer runtime. SDb converges in about the same number of iterations as KLD, but each iteration takes longer.
@@
Fig 9 compares the marginal densities estimated using MCMC with variational approximations from KLD, SDb and (p)BaM for some variables in Epi I and polypharmacy. For Epi I, all variational approaches match the MCMC results very closely except for $\zeta$, where SDb underestimated the marginal posterior variance more severely than BaM and KLD. For the higher-dimensional polypharmacy data, KLD matches MCMC results most closely, while SDb and pBaM tend to underestimate the marginal posterior variance although the mode was captured more accurately in some cases.
The results in this section are mixed, although KLD and (p)BaM tend to perform better than other methods. The superior performance of KLD may be related to the findings in Sections 2 and 3, which showed that the mean and variance underestimation for KLD is often less severe than FD or SD, leading to an overall higher accuracy. While SDb and BaM both minimize the batch approximated SD, BaM relies on closed form updates that lead to higher accuracy and faster convergence in low-dimensional problems, compared to SDb which is based on SGD. The advantages of SDb are more apparent for the higher-dimensional stochastic volatility models in Section 7.3, where (p)BaM fails to converge, and SDb surpasses KLD in overall accuracy and marginal mean and mode estimation. We hypothesize that SDb may perform better than KLD for skewed heavy-tailed posteriors, although this conjecture remains to be verified.
7.3 Stochastic volatility model
@@
where $\phi\in(0,1)$. To allow unconstrained updates, we apply the transformations, $\alpha=\log\sigma$ and $\psi=\text{\rm logit}(\phi)$. The set of local variables is $\theta_{L}=(b_{1},\dots,b_{n})^{\top}$ and global variables are $\theta_{G}=(\alpha,\lambda,\psi)^{\top}$. We consider the prior $\theta_{G}\sim{\text{N}}(0,\sigma_{0}^{2}I)$, where $\sigma_{0}^{2}=10$. For this model, $b_{i}$ is independent of $b_{j}$ given $\theta_{G}$ a posteriori if |ij|>1|i-j|>1. Thus, the Hessian of $\log h(\theta)$ has the same sparsity structure as $\Omega$ in the variational approximation. Both $\nabla_{\theta}\log(h(\theta))$ and $\nabla_{\theta}^{2}\log h(\theta)$ are derived in the supplement.
KLD FDr SDr FDb SDb
[…]
$\frac{\sigma}{\sigma^{*}}$ GBP 0.92 $\pm$ 0.05 0.68 $\pm$ 0.12 0.85 $\pm$ 0.19 0.48 $\pm$ 0.06 0.88 $\pm$ 0.05
DEM 0.95 $\pm$ 0.03 0.60 $\pm$ 0.08 0.99 $\pm$ 0.19 0.47 $\pm$ 0.04 0.91 $\pm$ 0.03
time GBP 9.2 (20) 18.9 (30) 8.3 (13) 1452.9 (30) 898.5 (14)
DEM 9.7 (19) 27.3 (30) 23.7 (25) 2748.0 (30) 1460.3 (13)
Table 7: Stochastic volatility model. Mean and standard deviation of normalized absolute difference in mode and mean and standard deviation ratio (best values highlighted in bold). Runtime is in seconds and number of iterations (in thousands) is given in brackets.
We analyze two datasets from Garch in the R package Ecdat. The first contains $n=1323$ observations of the weekday exchange rates of the U.S. Dollar against the British Pound (GBP) from 1 Aug 1980 to 28 Oct 1985. The second contains $n=1866$ observations of the weekday exchange rates for the U.S. Dollar against the German Deutschemark (DEM) from 2 Jan 1980 to 21 May 1987. For both datasets, the mean-corrected log-return series $\{y_{t}\}$ is derived from exchange rates $\{r_{t}\}$ using
@@
We set $B=10$ for FDb and SDb. We have tried various batchsizes for BaM, but the updates were severely ill-conditioned and BaM failed to converge. The challenge of inferring a full covariance matrix of dimension exceeding 1000 for BaM here is immense, further complicated by the high computational cost of matrix inversion. Although pBaM scales better to high dimensions than BaM, it also fails to converge for $B\in\{32,64,128\}$ and ranks $K\in\{8,16,32,64,128\}$. This may be because the posterior conditional independence structure of the stochastic volatility model is difficult to capture via a factor covariance matrix. The maximum number of iterations is set as $30,000$.
From Table 7, SDb provides the best approximations of the mean and mode, while KLD yields the most accurate estimates of the standard deviation. Note that SDr achieves a higher standard deviation ratio of $0.99$ for DEM, but does so with a much larger standard deviation of $0.19$, making KLD more reliable. In terms of runtime, KLD is the most efficient.
Fig 10 illustrates the impact of varying the batchsize for SDb in terms of convergence rate and approximation accuracy measured by $M^{*}$. Increasing the batchsize clearly leads to faster convergence and improved accuracy. The total runtime (shown in the legends of the first row) tends to decrease as fewer iterations are required for convergence. This suggests that larger batchsizes can enhance the stability and accuracy of SDb. Notably, the $M^{*}$ values of SDb exceed those of KLD even with a small $B=3$.
Figure 10: Stochastic volatility model. Lower bound averaged over every 1000 iterations and boxplots of $M^{*}$ values for different batchsizes for SDb.
Fig 11 compares the marginal posterior density estimates from MCMC, KLD and SDb ($B=10,100$) for some local variables and all global variables in DEM. SDb can capture the marginal posterior mode more accurately than KLD, especially for each of the global variables, but has a higher tendency to underestimate the posterior variance. Increasing the batchsize from 10 to 100 helps in reducing underestimation of the posterior variance.
Figure 11: Stochastic volatility model. Marginal density estimates for some local and global variables in DEM.